A 2.0 kg block starts from rest. It is pushed across a rough horizontal surface by a force F (magnitude of F = 8.0 N) directed at a 40° angle as shown. The block is found to accelerate along the surface at 2.0 m/s2.

Prove that there is a frictional force acting on the block

F-8.0 N 40° 40 2.0 kg
Answer #1

If friction does not present,

Fnetx = F*cos(40)

m*a = 8*cos(40)

a = 8*cos(40)/m

= 8*cos(40)/2

= 3.064 m/s^2

but the acceleration given is 2 m/s^2 < 3.064 m/s^2.

so, there is firstiobnal force , fk = m*(3.064 - 2)

= 2*(3.064 - 2)

= 2.128 N